# Particle transport and deposition in wall-sheared thermal turbulence

### v1

@@ Abstract
We studied the transport and deposition behavior of point particles in Rayleigh-Bénard convection cells subjected to Couette-type wall shear. Direct numerical simulations (DNS) are performed for Rayleigh number ($Ra$) in the range $10^{7}\leq Ra\leq 10^{9}$ with a fixed Prandtl number $Pr=0.71$, while the wall-shear Reynolds number ($Re_{w}$) is in the range $0\leq Re_{w}\leq 12000$. With the increase of $Re_{w}$, the large-scale rolls expanded horizontally, evolving into zonal flow in two-dimensional simulations or streamwise-oriented rolls in three-dimensional simulations. We observed that for particles with a small Stokes number ($St$), they either circulated within the large-scale rolls when buoyancy dominated or drifted near the walls when shear dominated. For medium $St$ particles, pronounced spatial inhomogeneity and preferential concentration were observed regardless of the prevailing flow state. For large $St$ particles, the turbulent flow structure had a minor influence on particles’ motion; although clustering still occurred, wall shear had a negligible influence compared to that for medium $St$ particles. We then presented the settling curves to quantify the particle deposition ratio on the walls. Our DNS results aligned well with previous theoretical predictions, which state that small $St$ particles settle with an exponential deposition ratio and large $St$ particles settle with a linear deposition ratio. For medium $St$ particles, where complex particles-turbulence interaction emerges, we developed a new model describing the settling process with an initial linear stage followed by a non-linear stage. Unknown parameters in our model can be determined either by fitting the settling curves or using empirical relations. Compared with DNS results, our model also accurately predicts the average residence time across a wide range of $St$ for various $Re_{w}$.
@@ 1 Introduction
Turbulence significantly affects the dynamics of particulate suspensions, as evidenced ubiquitously in nature and daily life (Guha, 2008; Toschi & Bodenschatz, 2009; Mathai et al., 2020; Brandt & Coletti, 2022). For example, aeolian processes can lead to the airborne movement or saltation of sand grains of different sizes under the influence of wind, triggering dust storms, especially under unstable thermal conditions (Zheng, 2009; Zhang, 2024). Another example is the use of heating, ventilation, and air conditioning (HVAC) systems, which can extend the suspension duration and spread of respiratory droplets, with this effect amplified by air currents in indoor environments (Mittal et al., 2020; Hedworth et al., 2021). In these scenarios, buoyancy forces generated by temperature differences drive turbulent flows. To study thermal turbulence, the Rayleigh-Bénard (RB) convection system serves as a canonical paradigm. This system involves a fluid layer heated from the bottom and cooled from the top (Ahlers et al., 2009; Lohse & Xia, 2010; Chillà & Schumacher, 2012; Wang et al., 2020a; Xia et al., 2023). The control parameters in the RB system include the Prandtl number ($Pr$), which describes a fluid’s thermophysical properties, and the Rayleigh number ($Ra$), which describes the relative strength of buoyancy forces versus thermal and viscous dissipation. In the RB turbulent convection, ubiquitous coherent structures include small-scale plumes and large-scale circulation (LSC). After detaching from the boundary layers, sheet-like plumes transform into mushroom-like plumes through mixing and clustering (Zhou et al., 2007). Through plume-vortex and plume-plume interactions, thermal plumes further self-organize into the LSC, which spans the whole convection cell (Xi et al., 2004).
@@
Due to these complex multi-scale coherent structures of thermal turbulence, efforts have been devoted to investigating the transport and deposition behavior of particles in thermal convection. A particularly interesting aspect is the spatial distribution and deposition rate of particles in thermal turbulence. Depending on the Stokes number ($St$), which describes the particle inertia relative to that of the fluid, the particles’ dynamic behavior can be classified into three categories. For particles with a small $St$, they are randomly distributed and behave like tracer particles; in the limit of infinitely small $St$, they exhibit an exponential deposition rate (Martin & Nokes, 1989). For particles with a large $St$, their motion is almost unaffected by the underlying thermal turbulence; in the limit of infinitely large $St$, they settle at a constant speed of terminal velocity $v_{t}$, and the deposition rate on the wall follows a linear law as derived from Stokes’ law. For particles with a medium $St$, they tend to cluster into band-like structures, and these structures are found to align with the vertical movement of plumes in thermal turbulence (Park et al., 2018). To predict the deposition rate for medium $St$, Patočka et al. (2020, 2022) developed a mathematical model that describes particle sedimentation as a stochastic process. Particles move from areas of intense convection to lower-velocity regions near the horizontal boundaries of the cell, with the possibility of escaping low-velocity regions without settling. In addition, the effects of thermal and mechanical coupling on the turbulence structure of the LSC and boundary layers, as well as particles’ motion, have received wide interest recently through simulations of two-way coupling (Oresta & Prosperetti, 2013; Park et al., 2018) or four-way coupling (Demou et al., 2022) between fluid and particles. Recent progress includes the works of Du & Yang (2022), Yang et al. (2022a, b), Sun et al. (2024), and Chen & Prosperetti (2024), which incorporate thermal conduction and radiation, accounting for the thermal backreaction on the flow and examining its implications on heat transfer efficiency and flow structure modulation.
@@
In addressing the complexities of natural and engineering fluid systems, such as atmospheric convection, oceanic currents, and indoor air ventilation, it is crucial to examine the interactions between vertical buoyancy and horizontal shear force (Hori et al., 2023). For example, Blass et al. (2020, 2021) incorporated a Couette-type shear into the RB system, where the top and bottom walls move in opposite directions at a constant speed. They found a transition in the flow dynamics from buoyancy-dominated to shear-dominated regimes as the strength of wall shear increased. In the buoyancy-dominated regime, the flow structure resembles canonical RB convection. In the shear-dominated regime, they observed the development of large meandering rolls. Yerragolam et al. (2022) then analyzed the spectra of convective flux and turbulent kinetic energy, thereby offering insights into the small-scale flow structures within the same convection cell. Further, Jin et al. (2022) reported enhanced interactions between the LSC and secondary flows over rough shearing surfaces, leading to an increased generation of thermal plumes. Xu et al. (2023) manipulated the movement of adiabatic sidewalls, thereby inducing vertical fluid motion that enhances heat transfer efficiency and can cause turbulent flows to relaminarize. In double diffusive convection, Li & Yang (2022) observed that even weak shear significantly alters the fingering morphology and transport properties of the system. Although recent advancements have shed light on the dynamics of single-phase and wall-sheared thermal turbulence, the complex interactions between a dispersed particulate phase and its surrounding fluid in wall-sheared thermal convection remain largely unexplored.
@@
In this work, we aim to investigate particle transport and deposition behavior within thermal turbulence under the influence of wall shearing. Motivated by studies of sand grains and pollutants dispersed in the air, we have chosen particles that are micrometers in size and much heavier than the surrounding fluid. Existing studies on particle-laden RB turbulence have largely focused on particles circulating within the LSC; here, we additionally consider scenarios where the flow structure changes under the influence of horizontal shear forces. The rest of this paper is organized as follows. In §$\S$ 2, we present numerical details for the direct numerical simulation (DNS) of wall-sheared thermal turbulence laden with point particles. In §$\S$ 3, we analyze particle transport behavior in the cell and their spatial distribution, followed by their deposition patterns and a mathematical model to describe particle deposition rates. In §$\S$ 4, the main findings of the present work are summarized.
@@ 2.1 Problem statement
We explore the dynamics of particles within a two-dimensional (2-D) convection cell of dimensions $L\times H$ (see figure 1a) and within a three-dimensional (3-D) convection cell of dimensions $L\times H\times W$ (see figure 1b). Here, $x$ represents the wall-shear direction, $y$ represents the wall-normal direction, and $z$ represents the spanwise direction. The top and bottom walls of the cell are maintained at constant temperatures of $T_{\emph{\text{cold}}}$ and $T_{\emph{\text{hot}}}$, respectively, while they move in opposite directions at a constant speed of $u_{w}$. The cell’s vertical walls are configured to be periodic. Our simulation protocol is as follows: we start the simulation of single-phase turbulent thermal convection, and particles are introduced into the turbulence once the flow reaches a statistically stationary state. Initially, the cell contains $N_{0}$ uniformly distributed stationary particles. We then advance the fluid flow and particle motion simultaneously, collecting statistics. We assume that particles adhere to the wall upon contact, ceasing further movement within the convection cell and no longer interacting with other deposited particles.
Figure 1: Schematic illustration of the wall-sheared convection cell and particles’ initial positions for: (a) the two-dimensional (2-D) simulations; (b) the three-dimensional (3-D) simulations.
@@ 2.2 Direct numerical simulation of thermal turbulence
Here, ${{\boldsymbol{u}}_{f}}$ is the fluid velocity, and $P$ and ${T_{f}}$ are the pressure and temperature of the fluid, respectively. The coefficients $\beta$, $\nu$ and $\alpha$ denote the thermal expansion coefficient, kinematic viscosity, and thermal diffusivity, respectively. Reference state variables are indicated by subscript zeros. $g$ represents gravitational acceleration, and $\hat{\boldsymbol{y}}$ denotes the unit vector parallel to gravity in the wall-normal direction.
@@
When an external wall shear is introduced, an additional control parameter of wall-shear Reynolds number ($Re_{w}=H{u_{w}}/\nu$) is needed. The competition between buoyancy and shear effects can be quantified by the bulk Richardson number as $Ri=Ra/(Re_{w}^{2}Pr)$ (Blass et al., 2020, 2021; Yerragolam et al., 2022).
We adopt the spectral element method implemented in the open-source Nek5000 solver (version v19.0) for our DNS. Further details on the spectral element method and the validation of the Nek5000 solver are available in Fischer (1997); Kooij et al. (2018). Previously, we verified the Nek5000 simulation results for wall-sheared thermal convection against those from an in-house GPU solver based on the lattice Boltzmann method (Xu et al., 2017; Xu & Li, 2023). The results from both the Nek5000 solver and our lattice Boltzmann solver demonstrated consistent results for wall-sheared thermal convection (Xu et al., 2023).
@@ 2.3 Kinematics of the particles
Here, $\boldsymbol{u}_{f}$ is the fluid velocity at the particle’s location, which is obtained by interpolating fluid velocity fields from the mesh to the particle locations. $m_{p}$ and $\boldsymbol{u}_{p}$ represent the mass and velocity of the particle, respectively. The particle response time is $\tau_{p}=\rho_{p}d_{p}^{2}/(18\mu_{f})$, where $d_{p}$ is the particle diameter. The drag correction factor $f(Re_{p})$ is a function of the particle Reynolds number $Re_{p}=d_{p}|\boldsymbol{u}_{f}-\boldsymbol{u}_{p}|/\nu_{f}$. We determined $f(R{e_{p}})$ in our simulations by dynamically calculating $Re_{p}$ for each particle and then applying the correction factor based on the formula $f(Re_{p})=1+0.15Re_{p}^{0.687}$ (Clift et al., 1978). In cases where $Re_{p}\ll 1$, we have $f(Re_{p})\approx 1$, indicating the validity of Stokes’ drag law. In this work, the 2-D cases can be approximated as a cross-section of the convection cell. Similar to Yang et al. (2022a, b), we considered the fluid layer thickness in the spanwise direction to be very thin and set it as $d_{p}$. On average, a specific particle will occupy a 3-D convection cell with thin sides equal to $d_{p}$. This approach allows for the coupling between fluid and particle features to be somewhat 3-D, while the flow characteristics tend to be generally 2-D. The same formulation for the particle response time has been used by Patočka et al. (2020) and Yang et al. (2022a, b) in their 2-D studies.
@@
Here, ${{\boldsymbol{x}}_{p}}$ denotes the position of the particle. For simplicity, we only consider the translational motion of isotropic spherical particles. To capture the anisotropic behavior of particles in turbulent flows, the orientation and rotation dynamics of inertial particles can be modeled as described by Challabotla et al. (2015) and Calzavarini et al. (2020). After non-dimensionalizing with the scales introduced above, we obtain
@@
Here, $\Lambda=\left({{\rho_{p}}-{\rho_{f}}}\right)/\left({{\rho_{p}}\beta{\Delta_{T}}}\right)$ denotes the buoyancy ratio, which describes the relative importance of particle buoyancy with respect to thermal buoyancy of the fluid. $S{t_{f}}={\tau_{p}}/{t_{f}}$ denotes the particle Stokes number, which describes the particle response time ${\tau_{p}}$ relative to the flow timescale quantified by the free-fall time ${t_{f}}=\sqrt{H/(g\beta{\Delta_{T}})}$. At small Stokes numbers, the particle response time $\tau_{p}$ is very small, making the term $f(Re_{p})(\mathbf{u}_{f}-\mathbf{u}_{p})/\tau_{p}$ large unless $\mathbf{u}_{p}$ is very close to $\mathbf{u}_{f}$. As a result, the particle velocity $\mathbf{u}_{p}$ quickly matches the fluid velocity $\mathbf{u}_{f}$. In this regime, the drag force term dominates the particle’s motion, and gravitational effects are minimal as particles follow the fluid flow. At large Stokes numbers, the particle response time $\tau_{p}$ is large, causing the term $f(Re_{p})(\mathbf{u}_{f}-\mathbf{u}_{p})/\tau_{p}$ to become small. Hence, the drag force term is negligible, and the particle motion is dominated by the gravitational term $(\rho_{p}-\rho_{f})\mathbf{g}/\rho_{p}$. In this regime, particles sediment with a velocity that increases linearly over time as the effect of gravity becomes more significant.
@@
For particles settling in a stationary fluid where $\mathbf{u}_{f}=0$, at terminal settling velocity $\mathbf{u}_{p}(t)=\mathbf{u}_{t}$, the particle’s acceleration is zero (i.e., $d\mathbf{u}_{p}/dt=0$). The speed of terminal velocity $v_{t}$ is then given by:

### v2

@@ Abstract
We studied the transport and deposition behaviour of point particles in Rayleigh-Bénard convection cells subjected to Couette-type wall shear. Direct numerical simulations (DNSs) are performed for Rayleigh number ($Ra$) in the range $10^{7}\leq Ra\leq 10^{9}$ with a fixed Prandtl number $Pr=0.71$, while the wall-shear Reynolds number ($Re_{w}$) is in the range $0\leq Re_{w}\leq 12000$. With the increase of $Re_{w}$, the large-scale rolls expanded horizontally, evolving into zonal flow in two-dimensional simulations or streamwise-oriented rolls in three-dimensional simulations. We observed that, for particles with a small Stokes number ($St$), they either circulated within the large-scale rolls when buoyancy dominated or drifted near the walls when shear dominated. For medium $St$ particles, pronounced spatial inhomogeneity and preferential concentration were observed regardless of the prevailing flow state. For large $St$ particles, the turbulent flow structure had a minor influence on the particles’ motion; although clustering still occurred, wall shear had a negligible influence compared with that for medium $St$ particles. We then presented the settling curves to quantify the particle deposition ratio on the walls. Our DNS results aligned well with previous theoretical predictions, which state that small $St$ particles settle with an exponential deposition ratio and large $St$ particles settle with a linear deposition ratio. For medium $St$ particles, where complex particle-turbulence interaction emerges, we developed a new model describing the settling process with an initial linear stage followed by a nonlinear stage. Unknown parameters in our model can be determined either by fitting the settling curves or using empirical relations. Compared with DNS results, our model also accurately predicts the average residence time across a wide range of $St$ for various $Re_{w}$. 111 This article may be downloaded for personal use only. Any other use requires prior permission of the author and Cambridge University Press. This article appeared in Xu et al., J. Fluid Mech. 999, A15 (2024) and may be found at https://doi.org/10.1017/jfm.2024.936.
@@ 1 Introduction
Turbulence significantly affects the dynamics of particulate suspensions, as evidenced ubiquitously in nature and daily life (Guha, 2008; Toschi & Bodenschatz, 2009; Mathai et al., 2020; Brandt & Coletti, 2022). For example, aeolian processes can lead to the airborne movement or saltation of sand grains of different sizes under the influence of wind, triggering dust storms, especially under unstable thermal conditions (Zheng, 2009; Zhang, 2024). Another example is the use of heating, ventilation and air conditioning systems, which can extend the suspension duration and spread of respiratory droplets, with this effect amplified by air currents in indoor environments (Mittal et al., 2020; Hedworth et al., 2021). In these scenarios, buoyancy forces generated by temperature differences drive turbulent flows. To study thermal turbulence, the Rayleigh-Bénard (RB) convection system serves as a canonical paradigm. This system involves a fluid layer heated from the bottom and cooled from the top (Ahlers et al., 2009; Lohse & Xia, 2010; Chillà & Schumacher, 2012; Wang et al., 2020a; Xia et al., 2023). The control parameters in the RB system include the Prandtl number ($Pr$), which describes a fluid’s thermophysical properties, and the Rayleigh number ($Ra$), which describes the relative strength of buoyancy forces vs thermal and viscous dissipation. In RB turbulent convection, ubiquitous coherent structures include small-scale plumes and large-scale circulation (LSC). After detaching from the boundary layers, sheet-like plumes transform into mushroom-like plumes through mixing and clustering (Zhou et al., 2007). Through plume-vortex and plume-plume interactions, thermal plumes further self-organize into the LSC, which spans the whole convection cell (Xi et al., 2004).
@@
Due to these complex multi-scale coherent structures of thermal turbulence, efforts have been devoted to investigating the transport and deposition behaviour of particles in thermal convection. A particularly interesting aspect is the spatial distribution and deposition rate of particles in thermal turbulence. Depending on the Stokes number ($St$), which describes the particle inertia relative to that of the fluid, the particles’ dynamic behaviour can be classified into three categories. For particles with a small $St$, they are randomly distributed and behave like tracer particles; in the limit of infinitely small $St$, they exhibit an exponential deposition rate (Martin & Nokes, 1989). For particles with a large $St$, their motion is almost unaffected by the underlying thermal turbulence; in the limit of infinitely large $St$, they settle at a constant speed of terminal velocity $v_{t}$, and the deposition rate on the wall follows a linear law as derived from Stokes’ law. For particles with a medium $St$, they tend to cluster into band-like structures, and these structures are found to align with the vertical movement of plumes in thermal turbulence (Park et al., 2018). To predict the deposition rate for medium $St$, Patočka et al. (2020, 2022) developed a mathematical model that describes particle sedimentation as a stochastic process. Particles move from areas of intense convection to lower-velocity regions near the horizontal boundaries of the cell, with the possibility of escaping low-velocity regions without settling. In addition, the effects of thermal and mechanical coupling on the turbulence structure of the LSC and boundary layers, as well as particles’ motion, have received wide interest recently through simulations of two-way coupling (Oresta & Prosperetti, 2013; Park et al., 2018) or four-way coupling (Demou et al., 2022) between the fluid and particles. Recent progress includes the works of Du & Yang (2022), Yang et al. (2022a, b), Sun et al. (2024) and Chen & Prosperetti (2024), which incorporate thermal conduction and radiation, accounting for the thermal backreaction on the flow and examining its implications for heat transfer efficiency and flow structure modulation.
@@
In addressing the complexities of natural and engineering fluid systems, such as atmospheric convection, oceanic currents and indoor air ventilation, it is crucial to examine the interactions between vertical buoyancy and the horizontal shear force (Hori et al., 2023). For example, Blass et al. (2020, 2021) incorporated a Couette-type shear into the RB system, where the top and bottom walls move in opposite directions at a constant speed. They found a transition in the flow dynamics from buoyancy-dominated to shear-dominated regimes as the strength of wall shear increased. In the buoyancy-dominated regime, the flow structure resembles canonical RB convection. In the shear-dominated regime, they observed the development of large meandering rolls. Yerragolam et al. (2022) then analysed the spectra of convective flux and turbulent kinetic energy, thereby offering insights into the small-scale flow structures within the same convection cell. Further, Jin et al. (2022) reported enhanced interactions between the LSC and secondary flows over rough shearing surfaces, leading to an increased generation of thermal plumes. Xu et al. (2023) manipulated the movement of adiabatic sidewalls, thereby inducing vertical fluid motion that enhances heat transfer efficiency and can cause turbulent flows to relaminarize. In double diffusive convection, Li & Yang (2022) observed that even weak shear significantly alters the fingering morphology and transport properties of the system. Although recent advancements have shed light on the dynamics of single-phase and wall-sheared thermal turbulence, the complex interactions between a dispersed particulate phase and its surrounding fluid in wall-sheared thermal convection remain largely unexplored.
@@
In this work, we aim to investigate particle transport and deposition behaviour within thermal turbulence under the influence of wall shearing. Motivated by studies of sand grains and pollutants dispersed in the air, we have chosen particles that are micrometres in size and much heavier than the surrounding fluid. Existing studies on particle-laden RB turbulence have largely focused on particles circulating within the LSC; here, we additionally consider scenarios where the flow structure changes under the influence of horizontal shear forces. The rest of this paper is organized as follows. In §$\S$ 2, we present numerical details for the direct numerical simulation (DNS) of wall-sheared thermal turbulence laden with point particles. In §$\S$ 3, we analyse the particles’ transport behaviour in the cell and their spatial distribution, followed by their deposition patterns and a mathematical model to describe particle deposition rates. In §$\S$ 4, the main findings of the present work are summarized.
@@ 2.1 Problem statement
We explore the dynamics of particles within a two-dimensional (2-D) convection cell of dimensions $L\times H$ (see figure 1a) and within a three-dimensional (3-D) convection cell of dimensions $L\times H\times W$ (see figure 1b). Here, $L$ is the length, $H$ is the height and $W$ is the width of the simulation domain, $x$ represents the wall-shear direction, $y$ represents the wall-normal direction and $z$ represents the spanwise direction. The top and bottom walls of the cell are maintained at constant temperatures of $T_{\emph{\text{cold}}}$ and $T_{\emph{\text{hot}}}$, respectively, while they move in opposite directions at a constant speed of $u_{w}$. The cell’s vertical walls are configured to be periodic. Our simulation protocol is as follows: we start the simulation of single-phase turbulent thermal convection, and particles are introduced into the turbulence once the flow reaches a statistically stationary state. Initially, the cell contains $N_{0}$ uniformly distributed stationary particles. We then advance the fluid flow and particle motion simultaneously, collecting statistics. We assume that particles adhere to the wall upon contact, ceasing further movement within the convection cell and no longer interacting with other deposited particles.
Figure 1: Schematic illustration of the wall-sheared convection cell and particles’ initial positions for: (a) the 2-D simulations; (b) the 3-D simulations.
@@ 2.2 Direct numerical simulation of thermal turbulence
Here, ${{\boldsymbol{u}}_{f}}$ is the fluid velocity, and $P$ and ${T_{f}}$ are the pressure and temperature of the fluid, respectively. The coefficients $\beta$, $\nu$ and $\alpha$ denote the thermal expansion coefficient, kinematic viscosity and thermal diffusivity, respectively. Reference state variables are indicated by subscript zeros. Also, $g$ represents gravitational acceleration, and $\hat{\boldsymbol{y}}$ denotes the unit vector parallel to gravity in the wall-normal direction.
@@
When an external wall shear is introduced, an additional control parameter, the wall-shear Reynolds number ($Re_{w}=H{u_{w}}/\nu$), is needed. The competition between buoyancy and shear effects can be quantified by the bulk Richardson number as $Ri=Ra/(Re_{w}^{2}Pr)$ (Blass et al., 2020, 2021; Yerragolam et al., 2022).
We adopt the spectral element method implemented in the open-source Nek5000 solver (version v19.0) for our DNS. Further details on the spectral element method and the validation of the Nek5000 solver are available in Fischer (1997) and Kooij et al. (2018). Previously, we verified the Nek5000 simulation results for wall-sheared thermal convection against those from an in-house GPU solver based on the lattice Boltzmann method (Xu et al., 2017; Xu & Li, 2023). The results from both the Nek5000 solver and our lattice Boltzmann solver demonstrated consistent results for wall-sheared thermal convection (Xu et al., 2023).
@@ 2.3 Kinematics of the particles
Here, $\boldsymbol{u}_{f}$ is the fluid velocity at the particle’s location, which is obtained by interpolating the fluid velocity fields from the mesh to the particle locations. The parameters $m_{p}$ and $\boldsymbol{u}_{p}$ represent the mass and velocity of the particle, respectively. The particle response time is $\tau_{p}=\rho_{p}d_{p}^{2}/(18\mu_{f})$, where $d_{p}$ is the particle diameter. The drag correction factor $f(Re_{p})$ is a function of the particle Reynolds number $Re_{p}=d_{p}|\boldsymbol{u}_{f}-\boldsymbol{u}_{p}|/\nu_{f}$. We determined $f(R{e_{p}})$ in our simulations by dynamically calculating $Re_{p}$ for each particle and then applying the correction factor based on the formula $f(Re_{p})=1+0.15Re_{p}^{0.687}$ (Clift et al., 1978). In cases where $Re_{p}\ll 1$, we have $f(Re_{p})\approx 1$, indicating the validity of Stokes’ drag law. In this work, the 2-D cases can be approximated as a cross-section of the convection cell. Similar to Yang et al. (2022a, b), we assume the fluid layer has a finite but very thin thickness $d_{p}$ in the spanwise direction. Both the fluid and particle motions are restricted in the spanwise direction, preventing movement along this axis. This restriction reduces the degrees of freedom, simplifying the dynamics. Consequently, while the overall flow is predominantly two-dimensional, the particles experience a 3-D environment due to the finite thickness of the fluid layer. The same formulation for the particle response time has been used by Patočka et al. (2020) and Yang et al. (2022a, b) in their 2-D studies.
@@
Here, ${{\boldsymbol{x}}_{p}}$ denotes the position of the particle. For simplicity, we only consider the translational motion of isotropic spherical particles. To capture the anisotropic behaviour of particles in turbulent flows, the orientation and rotation dynamics of inertial particles can be modelled as described by Challabotla et al. (2015) and Calzavarini et al. (2020). After non-dimensionalizing with the scales introduced above, we obtain
@@
Here, $\Lambda=\left({{\rho_{p}}-{\rho_{f}}}\right)/\left({{\rho_{p}}\beta{\Delta_{T}}}\right)$ denotes the buoyancy ratio, which describes the relative importance of particle buoyancy with respect to thermal buoyancy of the fluid and $S{t_{f}}={\tau_{p}}/{t_{f}}$ denotes the particle Stokes number, which describes the particle response time ${\tau_{p}}$ relative to the flow time scale quantified by the free-fall time ${t_{f}}=\sqrt{H/(g\beta{\Delta_{T}})}$. At small Stokes numbers, the particle response time $\tau_{p}$ is very small, making the term $f(Re_{p})(\mathbf{u}_{f}-\mathbf{u}_{p})/\tau_{p}$ large unless $\mathbf{u}_{p}$ is very close to $\mathbf{u}_{f}$. As a result, the particle velocity $\mathbf{u}_{p}$ quickly matches the fluid velocity $\mathbf{u}_{f}$. In this regime, the drag force term dominates the particles’ motion, and gravitational effects are minimal as particles follow the fluid flow. At large Stokes numbers, the particle response time $\tau_{p}$ is large, causing the term $f(Re_{p})(\mathbf{u}_{f}-\mathbf{u}_{p})/\tau_{p}$ to become small. Hence, the drag force term is negligible, and the particle motion is dominated by the gravitational term $(\rho_{p}-\rho_{f})\mathbf{g}/\rho_{p}$. In this regime, particles sediment with a velocity that increases linearly over time as the effect of gravity becomes more significant.
@@
For particles settling in a stationary fluid where $\mathbf{u}_{f}=0$, at terminal settling velocity $\mathbf{u}_{p}(t)=\mathbf{u}_{t}$, the particle’s acceleration is zero (i.e. $d\mathbf{u}_{p}/dt=0$). The speed of terminal velocity $v_{t}$ is then given by